OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

x+ (x+1) + ( x+2) = 267 3x+3=267 x=88 largest = 90

OpenStudy (anonymous):

idk i keep coming up with 89

OpenStudy (anonymous):

88 + 89 + 90 = 267, so largest is 90.

OpenStudy (anonymous):

x+(x+1)+(x+2)=267 3x+3=267 3x=267-3 3x=264 x=264/3 x=88 since the largest integer is x+2 you'll get 88+2=90